# CLIP-QDA: An Explainable Concept Bottleneck Model

### v2

@@ Abstract
In this paper, we introduce an explainable algorithm designed from a multi-modal foundation model, that performs fast and explainable image classification. Drawing inspiration from CLIP-based Concept Bottleneck Models (CBMs), our method creates a latent space where each neuron is linked to a specific word. Observing that this latent space can be modeled with simple distributions, we use a Mixture of Gaussians (MoG) formalism to enhance the interpretability of this latent space. Then, we introduce CLIP-QDA, a classifier that only uses statistical values to infer labels from the concepts. In addition, this formalism allows for both sample-wise and dataset-wise explanations. These explanations come from the inner design of our architecture, our work is part of a new family of greybox models, combining performances of opaque foundation models and the interpretability of transparent models. Our empirical findings show that in instances where the MoG assumption holds, CLIP-QDA achieves similar accuracy with state-of-the-art methods CBMs. Our explanations compete with existing XAI methods while being faster to compute.
@@ 1 Introduction
  • We propose two efficient and mathematically grounded XAI methods for model explanation, named CLIP-QDAlocal and CLIP-QDAglobal. These methods encompass both global and local assessments of why the model behaves. The global metric directly emanates from our Gaussian modeling approach, providing a comprehensive evaluation of CLIP-QDA’s performance. Additionally, our local metric draws inspiration from counterfactual analysis, furnishing insights into individual data points.
@@ 2.1 Contrastive Image Language Pre-training (CLIP)
CLIP (Radford et al., 2021) is a state-of-the-art model that can jointly process image and text inputs. The model was pre-trained on a large dataset of images and texts to learn to associate visual and textual concepts. Then, the capacity of CLIP to create a semantically rich encoding induced the creation of many emergent models in detection, few-shot learning, or image captioning.
@@ 2.2 CLIP-based Concept Bottleneck Models (CLIP-based CBMs)
The term Concept Bottleneck Model (CBM), as outlined in Koh et al. (2020), pertains to the implementation of a bottleneck reliant on human-specified concepts to execute a task, predominantly image classification. Consequently, the resultant algorithm inherently facilitates enhanced interpretability. While the term itself is relatively recent, it characterizes a lineage of methods that were employed in earlier research (Kumar et al., 2009; Lampert et al., 2009; Koh et al., 2020; Losch et al., 2019). However, despite the advantages offered by CBMs in terms of better understanding, early implementations encountered challenges stemming from the requirement for dedicated datasets. These datasets needed to encompass not only inputs and labels but also incorporate human-specified concepts for each sample.
@@ 2.3 Explainable AI
The literature shows a clear distinction between non-transparent (or blackbox) and transparent (or whitebox) models. Transparent models are characterized by their inherent explainability. These models can be readily explained due to their simplicity and easily interpretable features. Examples of such models include linear regression (Galton, 1886), logistic regression (McCullagh, 2019), and decision trees (Quinlan, 1986). In contrast, non-transparent models are inherently non-explainable. This category encompasses models that could have been explainable if they possessed simpler and more interpretable features (Galton, 1886; Quinlan, 1986), as well as models that inherently lack explainability, including deep neural networks. The distinction between these two types of models highlights the trade-off between model complexity and interpretability (Arrieta et al., 2020), with transparent models offering inherent explainability while non-transparent models allow for better performance, but require the use of additional techniques for explanations, named post-hoc methods. Commonly used post-hoc methods include visualization techniques, such as saliency maps, which highlight the influential features in an image that contribute to the model’s decision-making. Within this category of methods, notable examples include approaches such as Grad-CAM (Selvaraju et al., 2017), which generates activation maps by computing the gradients of the output labels. Sensitivity analysis (Cortez & Embrechts, 2011) represents another avenue, involving the analysis of model predictions by varying input data. Sample-wise explanation techniques are also used to explain the model from a local simplification of the model around a point of interest (Ribeiro et al., 2016; Plumb et al., 2018). Finally, feature relevance techniques aim at estimating the impact of each feature on the decision (Lundberg & Lee, 2017).
@@ 3.2.1 Notations and formalism
Finally, we define the classifier as a function $h_{{\bm{\omega}}^{h}}({\bm{z}})$ with parameters ${\bm{\omega}}^{h}$ that, given a vector ${\bm{z}}$, concepts the predicted class.
@@ 3.2.2 Gaussian modeling of CLIP’s latent space
To analyze the behavior of the CLIP latent space, we conduct a thorough examination of the distribution of CLIP scores. To elucidate our modeling approach, we suggest to visualize a large set of samples from $Z^{j}$ by observing the CLIP scores of an entire set of images taken from a toy example, which consists of images representing only cats and cars (see the Cats/Cars dataset in Section 4.1). In this instance, the concept denoted by “$j$” corresponds to “Pointy-eared”.
@@ 3.4 Explainable AI for Concept Bottleneck Models
Figure 4: The different source of explanation of CLIP-QDA. Global (CLIP-QDAglobal) vs. Local (CLIP-QDAlocal) explanations, offering insights into classifier behavior across the dataset and on individual samples, respectively. Post-hoc explanations with CLIP-SHAP and CLIP-LIME offering using traditional techniques.
@@ 3.4.1 Dataset-wise explanation with CLIP-QDAglobal
Note that the resulting value is no longer a distance since we lost the commutativity property. Examples of explanations based on this metric are given in Sections 4.5 and A.7. Also, an alternative way to produce dataset-wise explanations, oriented on example based explanations, are also available in Section A.4.
@@ 3.4.3 CLIP-LIME and CLIP-SHAP
4 Experiments
4.1 Experimental Setup
We evaluate our methods on ImageNet (Deng et al., 2009), PASCAL-Part (Donadello & Serafini, 2016), MIT Indoor Scenes dataset (Quattoni & Torralba, 2009), MonuMAI (Lamas et al., 2021) and Flowers102 (Nilsback & Zisserman, 2008). In addition to these well-established datasets, we introduce a custom dataset named Cats/Dogs/Cars dataset (Section A.2). To construct this dataset, we concatenated two widely recognized datasets, namely, the Kaggle Cats and Dogs Dataset (Cukierski, 2013) and the Standford Cars Dataset (Krause et al., 2013). Subsequently, we filtered the dataset to contain images of white and black animals and cars exclusively. This curation resulted in six distinct subsets: “Black Cars”, “Black Dogs”, “Black Cats”, “White Cars”, “White Dogs”, “White Cats”. The primary objective of this dataset is to facilitate experiments under conditions of substantial data bias, such as classifying white cats when the training data has only encountered white dogs and black cats. Specifically, we refer to two distinct scenarios (Table 1): one containing cats and cars of both colors (referred to as the unbiased setup) and the other one with only black cats and white cars (referred to as the biased setup). When none of theses scenarios are referred (for example in the $Del$ values of Table 4 and 5), the complete dataset is used. In its final form, the dataset comprises 6,436 images. The dataset will be shared after the reviewing process.
@@ 4.1 Experimental Setup
In addition to CLIP-QDA, we also test our method with linear regression as a classifier from CLIP scores, frequently named linear probe classifier. The objective here is to provide a comparison with existing CLIP-based CBMs (Yang et al., 2023; Yan et al., 2023a; b).
@@ 4.2 CBM Quantitative Evaluation Process
Assessing the quality of explanations has long been a challenge, given its subjective nature. While quantitative evaluation processes exist for evaluating performance in a general setup (Hedström et al., 2023), they may not be well-suited to the unique characteristics of CBMs. Consequently, we present a novel method designed to evaluate XAI solutions tailored to CBMs. This method comprises two metrics: the Deletion metric, gauging faithfulness to the model, and the Detection metric, quantifying faithfulness to the data.
@@
The deletion metric is adapted from the methodology introduced by Petsiuk et al. (2018). The procedure involves taking each sample from the test set and nullifying (i.e., setting to the average value of the score across classes) a certain number, $N_{null}$, of concepts. We nullify the concepts based on their importance order as determined by each explanation method. If nullifying the concepts leads to a significant decrease in performance, we consider it a successful selection of concepts that influenced the classifier’s decision. The intuition behind this idea is that if the concept is important, its absence will result in a loss of performance. Given the different values of accuracy obtained by nullifying $N_{null}$ concepts, we note $Acc(N_{null})$ we deduce the deletion score $Del$ by computing the area under the curve of $Acc(N_{null})$. The interest is to probe the ability to correctly order the important concept in the explanation:
@@
Here, $N_{max}$ represents a hyperparameter that defines the maximum number of deletions considered. The selection of this hyperparameter is critical: it balances between capturing the metric’s capacity to fit the model and ensuring that the computed inputs remain plausible. A default value of 9 is assigned to maintain this equilibrium. Note that we normalize the result by the maximum accuracy $Acc(0)$ to allow better comparisons among different classifiers.
@@
The detection metric evaluates the model’s ability to identify relevant concepts. For each sample $s$, a set of ground truth concepts $\mathcal{S}_{s}$ to detect is defined by an oracle. . Then, we construct a set of concepts $\mathcal{T}_{s}$ consisting of the top $|\mathcal{S}_{s}|$ concepts in the explanation to be tested. The resulting detection metric is the average ratio of agreements among all samples in the test dataset:
@@
4.3 Gaussian Prior Hypothesis
@@
Figure 6: Multivariate Q-Q plot for a Gaussian fit for the PASCAL-Part images that have the label “Aeroplane”.
Notably employing a less precise set of concepts can introduce disturbances, as evidenced in the observations. As indicated in Section 3.2.2, the ambiguity associated with certain concepts, such as “Multi-doored”, can lead to bimodal distributions (an airplane having one, multiple, or no doors). In Figure 7, we shows the histogram of the clip scores ${\bm{z}}$ of the images that have the class “aeroplane”. Compared to the histogram of less ambiguous cases (like in 3), we observe that the histogram presents anomalies, especially around the mean.
Figure 7: Histogram and Gaussian fitting of CLIP scores ${\bm{z}}$ for images that have the class “aeroplane” of the attribute “Multi-doored”.
Additionally, we performed a similar experiment with larger sets of concepts. We selected random subsets containing 10, 15, and 20 concepts from the PASCAL-Part set listed in Table 6. The outcomes are displayed in Figure 8. In this scenario, it becomes obvious that the Gaussian assumption is increasingly violated as the number of concepts grows. Indeed, as the number of concepts increases, the likelihood of encountering ambiguous concepts in each sample significantly rises, which undermines the feasibility of modeling the data as an unimodal Gaussian distribution.
@@
Figure 8: Multivariate Q-Q plot for a Gaussian fit for the PASCAL-Part images that have the label “Aeroplane”.
4.4 Assessing the Accuracy of XAI Methods
@@
In this section, we undertake a comparative evaluation of the performance of our CLIP-QDA classifier in contrast to other classifiers commonly employed in contemporary research. Precisely, we compare our method to the use of CLIP as a zero-shot classifier, by making predictions based on the class that produces the maximum CLIP score with the image embedding. We also use the Yan et al. (2023a) method, which consists of using a linear layer as the classifier of the CBM (see Figure 2). Additionally, we use Yang et al. (2023)’s method, involving the use of a class-concept matrix as the classifier of the CBM. Our assessment is based on the computation of accuracy metrics across three diverse datasets: PASCAL-Part, ImageNet, MIT scenes, and MonuMAI. A comprehensive description of the procedure for concept set acquisition is provided in Appendix A.3. All the CLIP-based CBMs are trained on the same sets of concepts denoted in Table 6. Furthermore, for contextualization, we include in the table the test accuracy of ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2021) trained in a supervised manner by using images as inputs. We also include Greybox XAI (Bennetot et al., 2022), X-NeSyL (Díaz-Rodríguez et al., 2022) and Morales Rodríguez et al. into our study, which offer explainability but require training data with annotated concepts. It is important to note that these methods, which require training from images, have significantly longer training times.
Results of these experiments are available in Table 2. The image encoder used in all our experiments based on CLIP-based CBMs is $ViT-L/14@336px$ provided from OpenAI’s public repository.
Table 2: Test set accuracy. On the top, are methods that require full training on images, and on the bottom, CBMs. Greybox XAI, X-NeSyL and Morales Rodríguez et al. requiring concept annotations, results on MIT scenes and ImageNet are not available.
@@
Our findings reveal that fine-tuning using CBM, either as a linear or a QDA probe, significantly improves performance, as evidenced by the increase in accuracy compared to using CLIP as a zero-shot classifier. This improvement is particularly pronounced on datasets dedicated to specialized tasks, such as MonuMAI. Additionally, CLIP CBMs tend to achieve performances comparable to networks trained from raw images, making these models appealing for image classification due to their reduced training cost in both time and resources, as well as their interpretability. Notably, CLIP-QDA demonstrates competitive performance, when compared to linear probe techniques. However, a slight decrease in performance is observed, especially in the PASCAL-Part dataset. This decline could be attributed to the use of a larger set of concepts specifically for this dataset, posing challenges to the Gaussian assumption and potentially affecting the applicability of our classifier.
@@
4.5 Assessing the Dataset-wise Interpretability of XAI Methods
4.5.1 Evaluation of CLIP-QDAlocal and CLIP-QDAglobal on the Cats/Dogs/Cars dataset
To evaluate the relevance of our explanations, we present a toy example from our Cats/Dogs/Cars dataset by constructing a CBM consisting of the concepts Table 6, plus the concepts “Black” and “White”. Next, we display the 10 most influential concepts according to our global metric (the top 10 concepts that have the highest Wasserstein distance) in both biased and unbiased setups. Results are presented in Figure 10.
(a) Unbiased setup
(b) Biased setup
Figure 10: Global explanation on subsets of Cats/Cars. Here, $c_{1}$=“Cat” and $c_{2}$=“Car”. Positive values indicate concepts that are more prevalent in cat images than car images, while negative values indicate concepts that are more common in car images compared to cat images. We display here only the top 10 concepts that have the highest Wasserstein distance (the concept “Black” is positioned $15^{th}$ in the unbiased setup).
We show an application of our local metric within the framework of our biased setup, as previously described. Subsequently, we feed an image of a white cat into our classifier (Figure 11). It is noteworthy that the image is misclassified as a car. Our local metric demonstrates sensitivity to the dataset’s color bias, corroborating the warning issued by the global explanation.
(a) Input image
Figure 11: Local explanation on subsets of Cats/Cars. On the right figure, the blue/red scale represents the scaled counterfactuals as in equation 6, the text in each box corresponds to the label predicted after the perturbation, followed to the concept changed to obtain its result (in parentheses).

### v3

@@ Abstract
In this paper, we introduce an explainable algorithm designed from a multi-modal foundation model, that performs fast and explainable image classification. Drawing inspiration from CLIP-based Concept Bottleneck Models (CBMs), our method creates a latent space where each neuron is linked to a specific word. Observing that this latent space can be modeled with simple distributions, we use a Mixture of Gaussians (MoG) formalism to enhance the interpretability of this latent space. Then, we introduce CLIP-QDA, a classifier that only uses statistical values to infer labels from the concepts. In addition, this formalism allows for both sample-wise and dataset-wise explanations. These explanations come from the inner design of our architecture, our work is part of a new family of greybox models, combining performances of opaque foundation models and the interpretability of transparent models. Our empirical findings show that in instances where the MoG assumption holds, CLIP-QDA achieves similar accuracy with state-of-the-art CBMs. Our explanations compete with existing XAI methods while being faster to compute.
@@ 1 Introduction
  • We propose two efficient and mathematically grounded XAI methods for model explanation, named CLIP-QDAlocal and CLIP-QDAglobal. These methods encompass both global and local assessments of how the model behaves. The global metric directly emanates from our Gaussian modeling approach, providing a comprehensive evaluation of CLIP-QDA’s performance. Additionally, our local metric draws inspiration from counterfactual analysis, furnishing insights into individual data points.
@@ 2.1 Contrastive Image Language Pre-training (CLIP)
CLIP (Radford et al., 2021) is a multi-modal model that can jointly process image and text inputs. The model was pre-trained on a large dataset of images and texts to learn to associate visual and textual concepts. Then, the capacity of CLIP to create a semantically rich encoding induced the creation of many emergent models in detection, few-shot learning, or image captioning.
@@ 2.2 CLIP-based Concept Bottleneck Models (CLIP-based CBMs)
The term Concept Bottleneck Model (CBM), as outlined in Koh et al. (2020), refers to to the implementation of a bottleneck reliant on human-specified concepts to execute a task, predominantly image classification. Consequently, the resultant algorithm inherently facilitates enhanced interpretability. While the term itself is relatively recent, it characterizes a lineage of methods that were employed in earlier research (Kumar et al., 2009; Lampert et al., 2009; Koh et al., 2020; Losch et al., 2019). However, despite the advantages offered by CBMs in terms of better understanding, early implementations encountered challenges stemming from the requirement for dedicated datasets. These datasets needed to encompass not only inputs and labels but also incorporate human-specified concepts for each sample.
@@ 2.3 Explainable AI
The literature shows a clear distinction between non-transparent (or blackbox) and transparent (or whitebox) models. Transparent models are characterized by their inherent explainability. These models can be readily explained due to their simplicity and easily interpretable features. Examples of such models include linear regression (Galton, 1886), logistic regression (McCullagh, 2019), and decision trees (Quinlan, 1986). In contrast, non-transparent models are inherently non-explainable. This category encompasses models that could have been explainable if they possessed simpler and more interpretable features (Galton, 1886; Quinlan, 1986), as well as models that inherently lack explainability, including deep neural networks. The distinction between these two types of models highlights the trade-off between model complexity and interpretability (Arrieta et al., 2020), with transparent models offering inherent explainability while non-transparent models allow for better performance but require the use of additional techniques for explanations, named post-hoc methods. Commonly used post-hoc methods include visualization techniques, such as saliency maps, which highlight the influential features in an image that contribute to the model’s decision-making. Within this category of methods, notable examples include approaches such as Grad-CAM (Selvaraju et al., 2017), which generates activation maps by computing the gradients of the output labels. Sensitivity analysis (Cortez & Embrechts, 2011) represents another avenue, involving the analysis of model predictions by varying input data. Sample-wise explanation techniques are also used to explain the model from a local simplification of the model around a point of interest (Ribeiro et al., 2016; Plumb et al., 2018). Finally, feature relevance techniques aim at estimating the impact of each feature on the decision (Lundberg & Lee, 2017).
@@ 3.2.1 Notations and formalism
Finally, we define the classifier as a function $h_{{\bm{\omega}}^{h}}({\bm{z}})$ with parameters ${\bm{\omega}}^{h}$ that, given a vector ${\bm{z}}$, outputs the predicted class.
@@ 3.2.2 Gaussian modeling of CLIP’s latent space
To analyze the behavior of the CLIP latent space, we conduct a thorough examination of the distribution of CLIP scores. To elucidate our modeling approach, we suggest to visualize a large set of samples from $Z^{j}$ by observing the CLIP scores of an entire set of images taken from a toy example, which consists of images representing only cats and cars (see the Cats/Dogs/Cars dataset in Section 4). In this instance, the concept denoted by “$j$” corresponds to “Pointy-eared”.
@@ 3.4 Explainable AI for Concept Bottleneck Models
Figure 4: The different source of explanation of CLIP-QDA. Global (CLIP-QDAglobal) vs. Local (CLIP-QDAlocal) explanations offer insights into classifier behavior across the dataset and on individual samples, respectively. Post-hoc explanations with CLIP-SHAP and CLIP-LIME use traditional XAI techniques.
@@ 3.4.1 Dataset-wise explanation with CLIP-QDAglobal
Note that the resulting value is no longer a distance since we lost the commutativity property. Examples of explanations based on this metric are given in Sections 5.4 and A.7. Also, an alternative way to produce dataset-wise explanations, oriented on example-based explanations, is also available in Section A.4.
@@ 3.4.3 CLIP-LIME and CLIP-SHAP
4 Experimental Setup
4.1 Datasets
We evaluate our methods on ImageNet (Deng et al., 2009), PASCAL-Part (Donadello & Serafini, 2016), MIT Indoor Scenes dataset (Quattoni & Torralba, 2009), MonuMAI (Lamas et al., 2021) and Flowers102 (Nilsback & Zisserman, 2008). In addition to these well-established datasets, we introduce a custom dataset named Cats/Dogs/Cars dataset (Section A.2). To construct this dataset, we concatenated two widely recognized datasets, namely, the Kaggle Cats and Dogs Dataset (Cukierski, 2013) and the Standford Cars Dataset (Krause et al., 2013). Subsequently, we filtered the dataset to contain images of white and black animals and cars exclusively. This curation resulted in six distinct subsets: “Black Cars”, “Black Dogs”, “Black Cats”, “White Cars”, “White Dogs”, “White Cats”. The primary objective of this dataset is to facilitate experiments under conditions of substantial data bias, such as classifying white cats when the training data has only encountered white dogs and black cats. Specifically, we refer to two distinct scenarios (Table 1): one containing cats and cars of both colors (referred to as the unbiased setup) and the other one with only black cats and white cars (referred to as the biased setup). When none of theses scenarios are referred to (for example in the $Del$ values of Table 4 and 5), the complete dataset is used. In its final form, the dataset comprises 6,436 images. Additional information is available in Section A.2.
@@ 4.1 Datasets
4.2 Baselines
Classifiers.
Here, we present the baseline used for comparing performance in terms of inference (i.e., accuracy) with other algorithms performing classification tasks. First, we evaluate various classifiers, which are represented as the trainable component illustrated in Figure 2. We denote the method proposed by Yan et al., which involves training a linear layer as a classifier from CLIP scores, as one of the baselines. As used notably in (Yan et al., 2023a; b), this approach is referred to as linear probe. Additionally, we adopt LaBo, introduced by Yang et al. (2023), which employs a class-concept matrix as the classifier of the CBMs. For all these methods, the same set of concepts is used, with a comprehensive description of the procedure for acquiring the concept set provided in Appendix A.3. We also incorporate previous CBMs such as Greybox XAI (Bennetot et al., 2022), X-NeSyL (Díaz-Rodríguez et al., 2022), and the method proposed by Morales Rodríguez et al. into our study, which provides explainability but requires additional training data related with annotated concepts. It is noteworthy that these methods, requiring training from images, entail significantly longer training times. Additionally, we assess methods widely used as classifiers that are not CBMs, including the use of CLIP as a zero-shot classifier. We also include ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2021) trained in a supervised manner using images as inputs. Further implementation details can be found in Section A.1.
XAI Methods.
To compare our proposed XAI methods (CLIP-QDAglobal, CLIP-QDAlocal, CLIP-LIME, and CLIP-SHAP) with existing works, we evaluated several methods, including those providing explanations at both the image and concept levels. At the concept level, we examined explanations provided by LaBo (Yang et al., 2023) and Yan et al. (Yan et al., 2023b). For image-based explanations, we assessed SHAP (Lundberg & Lee, 2017) and LIME (Ribeiro et al., 2016). Further implementation details can be found in Section A.1.
@@ 4.3 CBM Quantitative Evaluation Process
Assessing the quality of explanations has long been a challenge, given its subjective nature. While quantitative evaluation processes exist for evaluating performance in a general setup (Hedström et al., 2023), they may not be well-suited to the specific characteristics of CBMs. Consequently, we present a novel method designed to evaluate XAI solutions, tailored to CBMs. This method comprises two metrics: the Deletion metric, gauging faithfulness to the model, and the Detection metric, quantifying faithfulness to the data.
@@
The deletion metric is adapted from the methodology introduced by Petsiuk et al. (2018). The procedure involves taking each sample from the test set and nullifying (i.e., setting to the average value of the score across classes) a certain number, $N_{null}$, of concepts. We nullify the concepts based on their importance order as determined by each explanation method. If nullifying the concepts leads to a significant decrease in performance, we consider it a successful selection of concepts that influenced the classifier’s decision. The intuition behind this idea is that if the concept is important, its absence will result in a loss of performance. Given the different values of accuracy obtained by nullifying $N_{null}$ concepts, denoted by $Acc(N_{null})$, we deduce the deletion score $Del$ by computing the area under the curve of $Acc(N_{null})$. The interest is to probe the ability to correctly order the important concepts in the explanation:
@@
Here, $N_{max}$ represents a hyperparameter that defines the maximum number of deletions considered. The selection of this hyperparameter is critical: it balances between capturing the metric’s capacity to fit the model and ensuring that the computed inputs remain plausible. A default value of 9 is assigned to maintain this equilibrium. Note that we normalize the result by the maximum accuracy $Acc(0)$ to allow for better comparisons among different classifiers.
@@
The detection metric evaluates the model’s ability to identify relevant concepts. For each sample $s$, a set of ground truth concepts $\mathcal{S}_{s}$ to detect is defined by an oracle. Then, we construct a set of concepts $\mathcal{T}_{s}$ consisting of the top $|\mathcal{S}_{s}|$ concepts in the explanation to be tested. The resulting detection metric is the average ratio of agreements among all samples in the test dataset:
@@
5 Experiments
5.1 Assessing the Gaussian Prior Hypothesis
@@
Figure 6: Multivariate Q-Q plot illustrating the Gaussian fit for conditional modeling of the latent space, given a fixed label. The plot is specifically for PASCAL-Part images labeled as "Aeroplane".
Notably, employing a less precise set of concepts can introduce disturbances, as shown in the observations. As indicated in Section 3.2.2, the ambiguity associated with certain concepts, such as “Multi-doored”, can lead to bimodal distributions (an airplane having one, multiple, or no doors). In Figure 7, we show the histogram of the clip scores ${\bm{z}}$ of the images that have the class “aeroplane”. Compared to the histogram of less ambiguous cases (like in Figure 3), we observe that the histogram presents anomalies, especially around the mean.
Figure 7: Histogram and Gaussian fitting of CLIP scores ${\bm{z}}$ of the attribute “Multi-doored”, for images that have the class “aeroplane”.
Additionally, we performed a similar experiment with larger sets of concepts. We selected random subsets containing 10, 15, and 20 concepts from the PASCAL-Part set listed in Table 6. The outcomes are displayed in Figure 8. In this scenario, it becomes obvious that the Gaussian assumption is increasingly violated as the number of concepts grows. Indeed, as the number of concepts increases, the likelihood of encountering ambiguous concepts in each sample significantly increases, which undermines the feasibility of modeling the data as an unimodal Gaussian distribution.
@@
Figure 8: Multivariate Q-Q plot illustrating the Gaussian fit for conditional modeling of the latent space, given a fixed label. The plot is specifically for PASCAL-Part images labeled as "Aeroplane".
5.2 Assessing the Accuracy
@@
In this section, we undertake a comparative evaluation of the performance of our CLIP-QDA classifier in contrast to classifiers used in the baseline (see Section 4.2). Results of these experiments are available in Table 2. The image encoder used in all our experiments based on CLIP-based CBMs is $ViT-L/14@336px$ provided by OpenAI’s public repository.
Table 2: Test set accuracy. On the top are methods that require full training on images, and on the bottom, CBMs. Because Greybox XAI, X-NeSyL and Morales Rodríguez et al. require concept annotations, their results on MIT scenes and ImageNet are not available.
@@
Our findings reveal that fine-tuning using CBM, either as a linear or a QDA probe, significantly improves performance, as evidenced by the increase in accuracy compared to using CLIP as a zero-shot classifier. This improvement is particularly pronounced on datasets dedicated to specialized tasks, such as MonuMAI. Additionally, CLIP CBMs tend to achieve performances comparable to networks trained from raw images, making these models appealing for image classification due to their reduced training cost in both time and resources, as well as their interpretability. Notably, CLIP-QDA demonstrates competitive performance, when compared to linear probe techniques.
However, CLIP-QDA faces challenges in delivering competitive results for datasets like PASCAL-Part and ImageNet, which feature a significantly larger number of labels and concepts. The accuracy decline could be attributed to the use of a broader set of concepts tailored specifically for these datasets. This challenges the Gaussian assumption and potentially impacts the effectiveness of our classifier. Notably, there appears to be a correlation between the tests conducted on the different datasets and the observations depicted in Figure 8. For instance, while MonuMAI uses 20 concepts, MIT scenes use 25, PASCAL-Part employs 80, and ImageNet involves a staggering 5000 concepts.